OpenStudy (anonymous):

Solve: 21x - 24x^2 + 6x = 9 -24x^2 A. 1/2 B. 1/3 C. 3/2 D. 2/3

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

b, but i did it in my head so might be wrong

OpenStudy (anonymous):

I think... without writing anything down.

OpenStudy (anonymous):

Haha yeah what zbay said.

OpenStudy (anonymous):

Well with 2 of us getting the same it might be right!?!?!

OpenStudy (anonymous):

Haha I wouldn't trust my answer to such a degree ;) 21x - 24x^2 + 6x = 9 -24x^2 21x + 6x = 9 27x = 9 x = 9/27 = 1/3 Seems right. Only logical conclusion? Geniuses.